OpenStudy (sally6sa):

Solve 2x^2 -7x+2=0 by completing the square

OpenStudy (3mar):

Could I help you?

OpenStudy (welshfella):

sorry i have to go but Im sure 3mar can help you on this one

OpenStudy (abdumu16):

x = 7/4 - sqrt(33)/4 x = 7/4 + sqrt(33)/4

OpenStudy (3mar):

Thank you @welshfella But I think that @Sally6sa is not here! When @Sally6sa came back, (s)he would better to check that... I think this would help. Let me know what you got....

OpenStudy (sally6sa):

@3mar thank youuuuu

OpenStudy (3mar):

Don't mention it! @Sally6sa What matters is that you got the idea! Did you?